OpenStudy (anonymous):

derive. ln(5x^(7x) + tanh(x))

OpenStudy (anonymous):

Hmm Derive or Derivative (Differentiate)?

OpenStudy (anonymous):

sorry, um find the derivative.

OpenStudy (anonymous):

imstuck here. 7xln5+lntanhx

OpenStudy (anonymous):

i think i got it. is it 7x/5 + 1/sec^2hx

OpenStudy (anonymous):

I don't know derivative of tanhx \[\frac{1}{5x^{7x} + \tanh x}\times \left(\frac{d}{dx}5x^{7x} + \frac{d}{dx}\tanh x\right)\] I hope it helps you

OpenStudy (anonymous):

oh snap. the 5x^7x shouldnt have an x next to the 5. its just 5^7x